OpenStudy (anonymous):

Find a fundamental matrix for x'=(-3/4, 1/8, 1/2, -3/4)x. (I got (-3/4-r)^2-1/16=0, how do I find the roots from here?)

OpenStudy (anonymous):

Is that a 2x2 matrix? And if it is, should I interpret that as the values left to right (row at a time), or top to bottom (column at a time)?

OpenStudy (anonymous):

Yes, -3/4 and 1/8 on the left, 1/2 and -3/4 on the right.

OpenStudy (anonymous):

How do I find the roots?

OpenStudy (anonymous):

Do you mean the eigenvalues?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So then it's just a matter of solving for \(r\): \[\begin{align*}\left(-\frac{3}{4}-\lambda\right)^2-\frac{1}{16}&=0\\ \left(\frac{3}{4}+\lambda\right)^2&=\frac{1}{16}&&\text{factored out }(-1)^2\\ \frac{3}{4}+\lambda=\pm\frac{1}{4} \end{align*}\]

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

Well, you get \(\lambda_1=-\dfrac{1}{2}\) and \(\lambda_2=-1\), right? Next thing to do is find the associated eigenvectors.

OpenStudy (anonymous):

Got it.